Find the future values of these ordinary annuities. Compounding occurs once a year. Do not round intermediate calculations. Round your answers to the nearest cent.

$400 per year for 14 years at 4%.

$200 per year for 7 years at 2%.

$1,000 per year for 7 years at 0%.

Rework parts a, b, and c assuming they are annuities due. Future value of

$400 per year for 14 years at 4%:

Future value of $200 per year for 7 years at 2%:

Future value of $1,000 per year for 7 years at 0%:
Answer #1

Future value of annuity=Annuity[(1+rate)^time period-1]/rate

1.Future value=400[(1.04)^14-1]/0.04

=400*18.29191119

=$7316.76(Approx).

2.Future value=200[(1.02)^7-1]/0.02

=200*7.434283382

=$1486.86(Approx).

3.Future value=1000*7

=$7,000

Future value of annuity due=Future value of annuity*(1+rate)

1.Future value=7316.76*1.04

=$7609.44(Approx).

2.Future value=1486.86*1.02

=$1516.59(Approx).

3.Future value=1000*7

=$7,000